# Formation Control via Rotation Symmetry Constraints

### v1

@@ Abstract
This work introduces a distributed formation control strategy for multi-agent systems based only on rotation symmetry constraints. We propose a potential function that enforces inter-agent rotational symmetries, with its gradient defining the control law driving the agents toward a desired symmetric and planar configuration. We show that only (n1)(n-1) edges, the minimal connectivity requirement, are sufficient to implement the control strategy, where $n$ is the number of agents. We further augment the design to address the maneuvering problem, enabling the formation to undergo coordinated translations, rotations, and scalings along a predefined virtual trajectory. Simulation examples are provided to validate the effectiveness of the proposed method.
@@ I INTRODUCTION
The demand for distributed formation control schemes in multi-agent systems (MAS) has grown significantly in recent years, with applications ranging from UAV swarm coordination for mapping and surveillance [1], to satellite constellations coordination for efficient communication relays [2]. The role of a formation control law is to steer the agents into a desired spatial configuration in a distributed fashion. This is commonly achieved by imposing explicit geometric constraints between neighboring agents, such as distance-based schemes [3, 4], where inter-agent distances are fixed, or bearing-based schemes [5], where relative directions are maintained. In both cases, the desired target configuration is characterized using only local information shared between neighboring agents.
@@
In many formations the desired configuration exhibits spatial symmetries between the agents - rotations and/or reflections, often inherent from sensing coverage or communication requirements. The work [6] introduced an approach leveraging formation symmetries together with inter-agent distance constraints, drastically reducing the required inter-agent communication links as compared to other approaches. This motivates the question of whether it is possible to design a formation control scheme that relies solely on symmetry constraints.
@@
Graph theory provides the natural framework for modeling the decentralization, interaction topology, and geometric configuration of a MAS. Agents are represented as nodes (vertices), with their communication links as edges. A central challenge lies in balancing sparse information exchange while ensuring convergence to the desired configuration. To address this challenge, distance and bearing-based schemes leverage results from rigidity theory, relying on minimal infinitesimal rigidity (MIR) as a crucial architectural requirement to guarantee convergence to a desired shape [3, 5]. For distance-based approaches in $\mathbb{R}^{2}$, the MIR property requires at least (2n3)(2n-3) edges to uniquely determine the formation (up to translations, rotation, and so-called flip ambiguities), where $n$ is the number of agents.
@@
A relevant line of research builds on the concept of augmented Laplacian formulations such as complex-Laplacian approaches [7, 8, 9], where complex weights replace the standard scalar weights of the graph Laplacian, represent inter-agent interactions encoding rotations, translations, and scalings. Matrix-weighted Laplacian formulations, where the weights are structured matrices encoding inter-agent relations, offers another approach. This idea has appeared in several formation control settings: bearing-based control [5], where projection matrices enforce relative direction constraints; or more recent works [10], where the matrix weights play the same role as complex weights. These works demonstrate that matrix-weighted Laplacians form an ongoing field of research with multiple directions for achieving formation objectives.
In this work, we consider a group of $n$ agents required to arrange themselves into a geometric spatial pattern specified only by a set of inter-agent rotation symmetry constraints. We formalize these constraints in Euclidean space as point group isometries corresponding to rotations. Point groups can be classified by several families [11]. In this paper we restrict our study to planar cyclic rotations, enforced between designated agent pairs. Each agent has access to its state in $\mathbb{R}^{2}$, and may exchange this information only with neighboring agents, as determined by an undirected interaction network. The control objective is to design a distributed control law that drives the agents from any initial configuration to a desired configuration satisfying the required rotation symmetry constraints between the agents.
@@
This paper contributes with a foundation for solving the formation control problem solely under rotation constraints, providing a counterpart to the rigidity-based framework. We introduce a potential function that enforces rotational symmetries between neighboring agents, whose gradient yields a distributed control law driving the system toward the null-space of a symmetry-constraining matrix-weighted Laplacian. We show that (n1)(n-1) edges — the minimal connectivity requirement — is sufficient to guarantee convergence to the desired formation. To enhance flexibility, we present an augmentation of the control strategy, enabling the desired formation to be achieved while undergoing coordinated translations, rotations, and scalings according to a time-varying reference virtual trajectory, effectively addressing the formation maneuvering problem. Additionally, the effectiveness of the approach is also demonstrated through a numerical example extension in $\mathbb{R}^{3}$.
@@
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consists of two non-empty sets: $\mathcal{V}=\{1,...,n\}$ the set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ the set of edges. In this work, $\mathcal{G}$ is assumed to be undirected. The notation $ij\in\mathcal{E}$ indicates that agent $i$ can receive information from its neighboring agent $j$, and vice versa. Let $I_{n}\in\mathbb{R}^{n\times n}$ be the identity matrix, and $\mathds{1}_{n}\in\mathbb{R}^{n}$ be the all-one column vector of dimension $n$. Let $\otimes$ be the Kronecker product.
@@ II Symmetry in Graphs and Frameworks
The main focus of this work is to leverage the inherent symmetries of a formation to solve the formation control problem. In this direction, we first review notions from group theory and graph theory used to formally define symmetry.
@@ II-A Symmetry in Graphs
The order of a group is the number of its elements. A subset $B\subseteq\Gamma$ that is itself a group under $\circ$, is called a subgroup.
@@
That is, an automorphism preserves the adjacency structure of the graph. The identity permutation $\mathrm{id}$ is always an automorphism, and if $\psi$ is an automorphism, then so is its inverse $\psi^{-1}$. Moreover, the composition of two automorphisms is again an automorphism. These properties ensure that the set of all automorphisms of $\mathcal{G}$ forms a group under composition. This group is called the automorphism group of $\mathcal{G}$, denoted by $\mathrm{Aut}(\mathcal{G})$. One can express every permutation as a composition of disjoint cycles of the permutation. A cycle is a successive action of the permutation that sends a vertex back to itself, i.e., $i\to\psi(i)\to\psi(\psi(i))\to\cdots\to\psi^{k}(i)=i$, where $\psi^{k}=\underbrace{\psi\circ\cdots\circ\psi}_{k\text{ times }}$. Such a cycle is compactly written using the cycle notation, denoted by $(i\,\psi(i)\,\cdots\psi^{k-1}(i))$. The integer $k$ is the length of the cycle.
@@
Fig. 1 shows the cycle graph $C_{3}$. We can identify all the automorphisms of $\mathrm{Aut}(C_{3})$. Naturally, we can directly identify the identity permutation $\mathrm{id}$. Additionally, consider a counter-clockwise rotation by $120^{\circ}$ of $C_{3}$. This gives the automorphism (in cycle notation) $\psi_{1}=(1\,2\,3)$. We also have $\psi_{2}=\psi_{1}^{2}=\psi_{1}\circ\psi_{1}=(1\,3\,2)$, which can be interpreted geometrically as an additional rotation by $240^{\circ}$. Additional permutations can be found by considering reflections. Consider first the reflection about the vertical blue line, giving the permutation $\psi_{3}=(1)(2\,3)$. Similarly, the reflection about the red line yields $\psi_{4}=(3)(1\,2)$, and the reflection about the green line gives $\psi_{5}=(2)(1\,3)$. Thus, we have that $\mathrm{Aut}(C_{3})=\{\mathrm{id},\psi_{1},\ldots,\psi_{5}\}$ has 6 automorphisms.
@@
Note that we can choose $\Gamma=\{\mathrm{id},\psi_{1},\psi_{1}^{2}\}$, which corresponds to the subgroup of rotational automorphisms of $C_{3}$. In this case, $C_{3}$ can be considered as a $\Gamma$-symmetric graph, where any vertex can be mapped to any other under the rotation actions of $\Gamma$.
@@ II-B Symmetry in frameworks
The embedding of symmetric graphs in Euclidean space is of interest, especially for formation control problems. In this direction, we now consider symmetry of frameworks [12]. A framework in $\mathbb{R}^{2}$ is defined as the pair $(\mathcal{G},p)$, where $p:\mathcal{V}\rightarrow\mathbb{R}^{2}$ assigns each node in $\mathcal{G}$ a position in Euclidean space, used to represent the physical position of the agents in the network.
@@
In this work, we restrict our study to frameworks whose underlying graph $\mathcal{G}$ is the cycle graph $C_{n}$. By using the standard Schoenflies notation for point groups [11, 13], we consider the rotational symmetries described by the cyclic point group $\mathcal{C}_{n}$ of order $n\geq 1$. That is, $\mathcal{C}_{n}$ specifies the rotation symmetries that map the agents into one another under rotations about the origin.
We define $\Gamma_{r}\in\mathrm{Aut}(C_{n})$ to be the subgroup of rotational automorphisms of $C_{n}$, where each element $\tau(\gamma)$ is designed as a rotation about the origin by an angle $\theta=2\pi/n$. Then, in a planar setting ($\mathbb{R}^{2}$), $\tau(\Gamma_{r})$ coincides with the cyclic point group $\mathcal{C}_{n}$. We represent the elements $\tau(\gamma)$ by the standard rotation matrix
$R(\theta)=\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}\in SO(2).$
Thus, for any two vertices $u,v\in\mathcal{V}$ of a $\mathcal{C}_{n}$-symmetric framework, we denote by $\gamma_{uv}\in\Gamma_{r}$ the group element satisfying $\tau(\gamma_{uv})p_{u}=p_{v}$. Here, $\tau(\gamma_{uv})$ is represented by the rotation $R(\theta)$ and consequently, $\tau(\gamma_{vu})=R(\theta)^{T}=R(-\theta)$. In each case, the desired configuration satisfies condition (1), where agent positions are mapped into one another by the respective rotation in $\mathcal{C}_{n}$.
@@ III Symmetry-based Formation Control
We define an interaction graph $\mathcal{G}_{I}=(\mathcal{V},\mathcal{E}_{I})$ to specify which agents are able to exchange information. This graph is defined as a spanning tree subgraph of $C_{n}$, to ensure the minimal connectivity requirement between the agents in the MAS. For example, for a $\mathcal{C}_{4}$-symmetric formation in Fig. 2(a), the interaction edge set $\mathcal{E}_{I}=\{(1,2),(2,3),(3,4)\}$ satisfies the connectivity requirement.
Let $\Gamma_{r}\subseteq\mathrm{Aut}(C_{n})$ denote the subgroup of rotational automorphisms of $C_{n}$, and let $\mathcal{G}_{I}=(\mathcal{V},\mathcal{E}_{I})$ be the interaction graph defined as a spanning tree subgraph of $C_{n}$. The control objective is to design a distributed control law $u_{i}(t)$ for each agent such that, for every edge $uv\in\mathcal{E}_{I}$,
$\lim_{t\to\infty}\big\|p_{u}(t)-\tau(\gamma_{vu})\,p_{v}(t)\big\|=0.$ (3)
Here, $\gamma_{vu}\in\Gamma_{r}$ is the permutation mapping $v$ to $u$ and $\tau(\gamma_{vu})$ is the associated point group element representing a rotation predefined for that edge.
@@ III-A Symmetry-based Control Law
$F(p(t))=\frac{1}{2}\sum_{uv\in\mathcal{E}_{I}}\|p_{u}(t)-\tau(\gamma_{vu})p_{v}(t)\|^{2}.$ (4)
@@
Figure 3: The contribution $\tau(\gamma_{ji})p_{j}-p_{i}$ in the control law (6) drives $p_{i}$ to a symmetric position of $p_{j}$.
The closed-loop dynamics of each agent (6) has a straightforward geometric interpretation (see Fig. 3). The control law attempts to reduce the distance error between the term $\tau(\gamma_{ji})p_{j}$ and $p_{i}$.
We now focus on the closed-loop dynamics in state-space
@@
$Q\in\mathbb{R}^{2n\times 2n}$ is the resulting symmetry-constraining matrix-weighted Laplacian for graph $\mathcal{G}_{I}$, with the block entries
$Q]_{uv}=\begin{cases}d(u)I_{2},&u=v,\,u\in\mathcal{V}\\ -\tau(\gamma_{vu}),&vu\in\mathcal{E}_{I}\\ 0,&\text{o.w.}\end{cases},$
@@
Consider the $\mathcal{C}_{4}$-symmetric framework seen in Fig. 2(a), with the choice of the edge set $\mathcal{E}_{I}=\{(1,2),(2,3),(3,4)\}$. Then, the group actions in $\tau(\Gamma_{r})$ are rotations about the origin by $2\pi/n=\pi/2$, and the corresponding matrix $Q\in\mathbb{R}^{8\times 8}$ can be expressed as
$Q=\begin{bmatrix}I_{2}&-R(\frac{\pi}{2})^{T}&0&0\\ -R(\frac{\pi}{2})&2I_{2}&-R(\frac{\pi}{2})^{T}&0\\ 0&-R(\frac{\pi}{2})&2I_{2}&-R(\frac{\pi}{2})^{T}\\ 0&0&-R(\frac{\pi}{2})&I_{2}\end{bmatrix}$.
@@
Let $Q$ be the symmetry-constraining matrix-weighted Laplacian associated with the spanning tree graph $\mathcal{G}_{I}$. Then:
  1. i)
    $Q$ is positive semi-definite (PSD).
  2. ii)
    $Q$ has a nontrivial null-space,
    $\mathrm{Null}(Q)=\{p\in\mathbb{R}^{2n}|E(\Gamma_{r})^{T}p=0\},$
    corresponding to the set of $\mathcal{C}_{n}$-symmetric configurations.
  3. iii)
    The rank of $Q$ is 2n22n-2, and $\dim\mathrm{Null}(Q)=2$.
@@
(i) Since $Q=E(\Gamma_{r})E(\Gamma_{r})^{T}$, for any $p\in\mathbb{R}^{2n}$:
$p^{T}Qp=p^{T}E(\Gamma_{r})E(\Gamma_{r})^{T}p=\|E(\Gamma_{r})^{T}p\|\geq 0.$
Hence, $Q$ is PSD.
(ii) Without loss of generality, assume the nodes are labeled such that each edge in $\mathcal{G}_{I}$ is of the form $i(i+1)$. By construction, $E(\Gamma_{r})^{T}p$ stacks the edge errors $r_{i}=p_{i}-\tau(\gamma_{(i(i+1))})^{T}\,p_{(i+1)}\in\mathbb{R}^{2}$, and
$E(\Gamma_{r})^{\top}p=\begin{bmatrix}r_{1}\\ \vdots\\ r_{|\mathcal{E}_{I}|}\end{bmatrix}\in\mathbb{R}^{2|\mathcal{E}_{I}|}.$
We have $E(\Gamma_{r})^{T}p=0$ iff every edge satisfies $p_{u}-\tau(\gamma_{vu})p_{v}=0$. Therefore, by Definition 4, $\mathrm{Null}(Q)$ coincides with the respective set of $\mathcal{C}_{n}$-symmetric configurations.
(iii) Under the assumption that $\mathcal{G}_{I}$ is a spanning tree, for each node $i$ define the matrix $S_{i}\in SO(2)$, with $S_{1}=I_{2}$, and let
$S_{i+1}=\tau(\gamma_{i(i+1)})S_{i}\in SO(2)$ (8)
be the ordered product of edge rotations along the unique path from $1$ to $j$. Define $p_{1}=q\in\mathbb{R}^{2}$. Then, if $E(\Gamma_{r})^{T}p=0$, by substitution we have
$p_{i}\;=\;S_{i}\,q,\qquad\forall i\in\{1,\dots,n\}.$
such that
$E(\Gamma_{r})^{T}\begin{bmatrix}S_{1}q\\ S_{2}q\\ \vdots\\ S_{n}q\end{bmatrix}=0.$
Note that $q\in\mathbb{R}^{2}$ has two degrees of freedom.
Hence any vector in $\mathrm{Null}(E(\Gamma_{r})^{T})=\mathrm{Null}(Q)=\mathrm{IM}(V_{0})$ where
$V_{0}=\begin{bmatrix}S_{1}e_{1}&S_{1}e_{2}\\ S_{2}e_{1}&S_{2}e_{2}\\ \vdots&\vdots\\ S_{n}e_{1}&S_{n}e_{2}\end{bmatrix}$, (9)
with $e_{1}=\begin{bmatrix}1&0\end{bmatrix}^{T}$ and $e_{2}=\begin{bmatrix}0&1\end{bmatrix}^{T}$. Therefore, $\dim\mathrm{Null}(Q)=2$, and
$\mathrm{rank}(Q)=2n-\dim\mathrm{Null}(Q)=2n-2.$
We now examine the dynamics of the closed-loop system (7) to show that the proposed distributed control law drives the agents from any initial condition to the desired symmetric configuration. We next show the explicit solution of (7), showing that the limit configuration corresponds to the orthogonal projection of the initial state onto the subspace of $\mathcal{C}_{n}$-symmetric formations.
@@
Consider a MAS consisting of $n$ integrator agents (2), whose interaction topology is defined by a spanning tree graph $\mathcal{G}_{I}$, and let
@@
where $V_{0}$ is given in (9).
Furthermore, the steady-state of each agent is given by
$\lim_{t\rightarrow\infty}p_{i}(t)=\frac{1}{n}S_{i}\sum_{k=1}^{n}S_{k}^{T}p_{k}(0),$ (11)
with $S_{i}$ defined as in (8).

### v2

@@ Abstract
This work introduces a distributed formation control strategy for multi-agent systems based solely on rotation symmetry constraints. We propose a potential function that enforces inter-agent rotational symmetries, whose gradient defines a control law that drives the agents toward a desired planar symmetric configuration. We show that only n1n-1 edges (the minimal connectivity requirement) are sufficient to implement the strategy, where $n$ is the number of agents. We further augment the design to address the maneuvering problem, enabling the formation to undergo coordinated translations, rotations, and scaling along a predefined virtual trajectory. Simulation examples are provided to validate the effectiveness of the proposed method.
@@ I INTRODUCTION
The demand for distributed formation control schemes in multi-agent systems (MAS) has grown significantly in recent years, with applications ranging from UAV swarm coordination for mapping and surveillance [1] to satellite constellation coordination for efficient communication relays [2]. The role of a formation control scheme is to provide a control law that steers the agents into a desired spatial configuration in a distributed fashion. This is commonly achieved by imposing explicit geometric constraints between neighboring agents. For example, distance-based schemes [3, 4] fix inter-agent distances, while bearing-based schemes [5] fix relative directions. In both cases, the desired target configuration is characterized using only local information shared between neighboring agents.
@@
In many formations, the desired configuration exhibits spatial symmetries between agents such as rotations and/or reflections, often inherent to sensing coverage or communication requirements. The work [6] introduced an approach that leverages formation symmetries together with inter-agent distance constraints, drastically reducing the required inter-agent communication links compared with traditional distance-based approaches. This motivates the question of whether it is possible to design a formation control scheme that relies solely on symmetry constraints.
@@
Graph theory provides a natural framework for modeling the decentralization, interaction topology, and geometric configuration of a MAS. Agents are represented as nodes (vertices), with communication links as edges. A central challenge lies in balancing sparse information exchange while ensuring convergence to the desired configuration. To address this challenge, distance-based and bearing-based schemes leverage results from rigidity theory, relying on minimal infinitesimal rigidity (MIR) as a crucial architectural requirement to guarantee convergence to a desired shape [3, 5]. For distance-based approaches in $\mathbb{R}^{2}$, the MIR property requires at least 2n32n-3 edges to uniquely determine the formation (up to translations, rotation, and so-called flip ambiguities), where $n$ is the number of agents.
@@
A closely related line of research builds on augmented Laplacian formulations, such as complex-Laplacian approaches [7, 8, 9], where complex weights replace the standard scalar weights of the graph Laplacian and encode inter-agent rotations, translations, and scalings. Matrix-weighted Laplacians provide a related framework, where structured matrix weights encode inter-agent geometric relations. This idea has appeared in several formation control settings: bearing-based control [5], where projection matrices enforce relative direction constraints, and more recent works [10], where matrix weights play the same role as complex weights. Together, these works highlight the potential of matrix-weighted Laplacians for achieving formation control objectives.
@@
Motivated by this line of research, we build on [6] to study formations defined solely through symmetric relations. Specifically, consider a group of $n$ agents whose desired configuration is characterized by prescribed rotational symmetries. These constraints are modeled in Euclidean space via cyclic point-group isometries [11], enforced between designated agent pairs. Assume that each agent has access to its own state and may exchange this information only with neighboring agents, as determined by an undirected interaction graph. The control objective is to design a distributed control law that drives the agents from any initial state to a desired configuration satisfying the required rotation symmetry constraints. This is achieved by introducing a potential function that enforces rotational symmetries between neighboring agents, whose gradient yields a distributed control law that drives the system toward the null space of a symmetry-constraining matrix-weighted Laplacian. We show that n1n-1 edges (the minimal connectivity requirement) are sufficient to guarantee convergence to the desired formation. To enhance flexibility, we present an augmentation of the control strategy, enabling the desired formation to be achieved while undergoing coordinated translations, rotations, and scalings according to a time-varying reference virtual trajectory, thereby addressing the formation maneuvering problem. Additionally, the effectiveness of the approach is demonstrated through a numerical example extension in $\mathbb{R}^{3}$.
@@
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consists of two non-empty sets: $\mathcal{V}=\{1,...,n\}$, the set of nodes, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$, the set of edges. In this work, $\mathcal{G}$ is assumed to be undirected. The notation $ij\in\mathcal{E}$ indicates that agent $i$ can receive information from neighboring agent $j$, and vice versa. Let $I_{n}\in\mathbb{R}^{n\times n}$ be the identity matrix, and let $\mathds{1}_{n}\in\mathbb{R}^{n}$ be the all-one column vector of dimension $n$. Let $\otimes$ denote the Kronecker product.
@@ II Symmetry in Graphs and Frameworks
The main focus of this work is to leverage the inherent symmetries of a desired configuration to solve the formation control problem. In this direction, we first review notions from group theory and graph theory used to formally define symmetry.
@@ II-A Symmetry in Graphs
The order of a group is the number of its elements. A subset $B\subseteq\Gamma$ that is itself a group under $\circ$ is called a subgroup.
@@
That is, an automorphism preserves the adjacency structure of the graph. The identity permutation $\mathrm{id}$ is always an automorphism, and if $\psi$ is an automorphism, then so is its inverse $\psi^{-1}$. Moreover, the composition of two automorphisms is again an automorphism. These properties ensure that the set of all automorphisms of $\mathcal{G}$ forms a group under composition. This group is called the automorphism group of $\mathcal{G}$, denoted by $\mathrm{Aut}(\mathcal{G})$. One can express every permutation as a composition of disjoint cycles. A cycle is a successive action of the permutation that sends a vertex back to itself, i.e., $i\to\psi(i)\to\psi(\psi(i))\to\cdots\to\psi^{k}(i)=i$, where $\psi^{k}=\underbrace{\psi\circ\cdots\circ\psi}_{k\text{ times }}$. Such a cycle is compactly written using the cycle notation, denoted by $(i\,\psi(i)\,\cdots\psi^{k-1}(i))$. The integer $k$ is the length of the cycle.
@@
Fig. 1 shows the cycle graph $C_{3}$. We can identify all the automorphisms of $\mathrm{Aut}(C_{3})$. We first identify the identity permutation $\mathrm{id}$. Additionally, consider a counter-clockwise rotation by $120^{\circ}$ of $C_{3}$. This gives the automorphism (in cycle notation) $\psi_{1}=(1\,2\,3)$. We also have $\psi_{2}=\psi_{1}^{2}=\psi_{1}\circ\psi_{1}=(1\,3\,2)$, which can be interpreted geometrically as a rotation by $240^{\circ}$. Additional permutations can be found by considering reflections. Consider first the reflection about the vertical blue line, giving the permutation $\psi_{3}=(1)(2\,3)$. Similarly, the reflection about the red line yields $\psi_{4}=(3)(1\,2)$, and the reflection about the green line gives $\psi_{5}=(2)(1\,3)$. Thus, $\mathrm{Aut}(C_{3})=\{\mathrm{id},\psi_{1},\ldots,\psi_{5}\}$ has 6 automorphisms.
@@
Note that the choice $\Gamma=\{\mathrm{id},\psi_{1},\psi_{1}^{2}\}$ corresponds to the subgroup of rotational automorphisms of $C_{3}$. In this case, $C_{3}$ can be considered as a $\Gamma$-symmetric graph, where any vertex can be mapped to any other under the rotation actions of $\Gamma$.
@@ II-B Symmetry in frameworks
The embedding of symmetric graphs in Euclidean space is of interest, especially for formation control problems. In this direction, we now consider symmetry of frameworks [12]. A framework in $\mathbb{R}^{2}$ is defined as the pair $(\mathcal{G},p)$, where $p:\mathcal{V}\rightarrow\mathbb{R}^{2}$ assigns each node in $\mathcal{G}$ a position in Euclidean space, which represents the physical position of the agents in the network.
@@
In this work, we restrict our study to frameworks whose underlying graph $\mathcal{G}$ is the cycle graph $C_{n}$. Using the standard Schoenflies notation for point groups [11, 13], we consider the rotational symmetries described by the cyclic point group $\mathcal{C}_{n}$ of order $n\geq 1$. That is, $\mathcal{C}_{n}$ specifies the rotation symmetries that map agents to one another under rotations about the origin.
We define $\Gamma_{r}\in\mathrm{Aut}(C_{n})$ as the subgroup of rotational automorphisms of $C_{n}$, where each element $\tau(\gamma)$ is represented as a rotation about the origin by an angle $\theta=2\pi/n$. Then, in a planar setting (i.e., in $\mathbb{R}^{2}$), $\tau(\Gamma_{r})$ coincides with the cyclic point group $\mathcal{C}_{n}$. We represent the elements $\tau(\gamma)$ by the standard rotation matrix $R(\theta)\in SO(2).$ Thus, for any two vertices $i,j\in\mathcal{V}$ of a $\mathcal{C}_{n}$-symmetric framework, we denote by $\gamma_{ij}\in\Gamma_{r}$ the group element satisfying $\tau(\gamma_{ij})p_{i}=p_{j}$ with $\tau(\gamma_{ij})=R(\theta)$ and consequently, $\tau(\gamma_{ji})=R(\theta)^{T}$. In each case, the desired configuration satisfies condition (1), where agent positions are mapped to one another by the corresponding rotation in $\mathcal{C}_{n}$.
@@ III Symmetry-based Formation Control
We define an interaction graph $\mathcal{G}_{I}=(\mathcal{V},\mathcal{E}_{I})$ to specify which agents are able to exchange information. This graph is defined as a spanning tree subgraph of $C_{n}$, to ensure the minimal connectivity requirement between agents in the MAS. For example, for a $\mathcal{C}_{4}$-symmetric formation in Fig. 2(a), the interaction edge set $\mathcal{E}_{I}=\{12,23,34\}$ satisfies the connectivity requirement.
Let $\Gamma_{r}\subseteq\mathrm{Aut}(C_{n})$ denote the subgroup of rotational automorphisms of $C_{n}$, and let $\mathcal{G}_{I}=(\mathcal{V},\mathcal{E}_{I})$ be the interaction graph defined as a spanning tree subgraph of $C_{n}$. The control objective is to design a distributed control law $u_{i}(t)$ for each agent $i$ such that, for every edge $ij\in\mathcal{E}_{I}$, we have
$\lim_{t\to\infty}\big\|p_{i}(t)-\tau(\gamma_{ji})\,p_{j}(t)\big\|=0.$ (3)
Here, $\gamma_{ji}\in\Gamma_{r}$ is the permutation mapping $j$ to $i$ and $\tau(\gamma_{ji})$ is the associated point group element representing a rotation predefined for that edge.
@@ III-A Symmetry-based Control Law
$F(p(t))=\frac{1}{2}\sum_{ij\in\mathcal{E}_{I}}\|p_{i}(t)-\tau(\gamma_{ji})p_{j}(t)\|^{2}.$ (4)
@@
The closed-loop dynamics of each agent (6) has a straightforward geometric interpretation. The control law attempts to reduce the error vector $p_{i}-\tau(\gamma_{ji})p_{j}$, i.e., it attempts to align the state of agent $i$ with the rotated state of its neighbor $j$ according to the prescribed symmetry action.
We now focus on the closed-loop dynamics in state-space, which has the form
@@
where $Q\in\mathbb{R}^{2n\times 2n}$ is the resulting symmetry-constraining matrix-weighted Laplacian for graph $\mathcal{G}_{I}$, with the block entries
$Q]_{ij}=\begin{cases}d(i)I_{2},&i=j,\,i\in\mathcal{V}\\ -\tau(\gamma_{ji}),&ij\in\mathcal{E}_{I}\\ 0,&\text{o.w.}\end{cases},$
@@
Consider the $\mathcal{C}_{4}$-symmetric framework seen in Fig. 2(a), with the choice of the edge set $\mathcal{E}_{I}=\{12,23,34\}$. Then, the group actions in $\tau(\Gamma_{r})$ are rotations about the origin by $2\pi/n=\pi/2$, and the corresponding matrix $Q\in\mathbb{R}^{8\times 8}$ can be expressed as
$Q=\left[\begin{smallmatrix}I_{2}&-R(\frac{\pi}{2})^{T}&0&0\\ -R(\frac{\pi}{2})&2I_{2}&-R(\frac{\pi}{2})^{T}&0\\ 0&-R(\frac{\pi}{2})&2I_{2}&-R(\frac{\pi}{2})^{T}\\ 0&0&-R(\frac{\pi}{2})&I_{2}\end{smallmatrix}\right].$
@@
Let $Q$ be the symmetry-constraining matrix-weighted Laplacian associated with $\mathcal{G}_{I}$. Then, (i) $Q$ is positive semi-definite (PSD). (ii) $Q$ has a nontrivial null-space, $\mathrm{Null}(Q)=\{p\in\mathbb{R}^{2n}|E(\Gamma_{r})^{T}p=0\}$, corresponding to the set of $\mathcal{C}_{n}$-symmetric configurations, and (iii) the rank of $Q$ is 2n22n-2, and $\dim\mathrm{Null}(Q)=2$.
@@
(ii) Since $Q=E(\Gamma_{r})E(\Gamma_{r})^{T}$, for any $p\in\mathbb{R}^{2n}$ we have $p^{T}Qp=p^{T}E(\Gamma_{r})E(\Gamma_{r})^{T}p=\|E(\Gamma_{r})^{T}p\|^{2}\geq 0.$ Hence, $Q$ is PSD. ($ii$) Without loss of generality, assume the nodes are labeled such that each edge in $\mathcal{G}_{I}$ is of the form $i(i+1)$. By construction, $E(\Gamma_{r})^{T}p=\begin{bmatrix}r_{1}^{T}&\cdots&r_{|\mathcal{E}_{I}|}^{T}\end{bmatrix}^{T}\in\mathbb{R}^{2|\mathcal{E}_{I}|}$ stacks the edge errors $r_{i}=p_{i}-\tau(\gamma_{(i(i+1))})^{T}\,p_{(i+1)}\in\mathbb{R}^{2}$. Then, $E(\Gamma_{r})^{T}p=0$ iff every edge satisfies $p_{i}-\tau(\gamma_{ji})p_{j}=0$. Therefore, by Definition 4, $\mathrm{Null}(Q)$ coincides with the respective set of $\mathcal{C}_{n}$-symmetric configurations. (iiiiii) Define $S_{1}=I_{2}$ and, for each node $j$, let $S_{j}\in SO(2)$ denote the ordered product of edge rotations along the unique path from node $1$ to node $j$, so that along each edge $i(i+1)$ we have $S_{i+1}=\tau(\gamma_{i(i+1)})S_{i}\in SO(2)$. Let $p_{1}=q\in\mathbb{R}^{2}$. Then, if $E(\Gamma_{r})^{T}p=0$, by substitution we have $p_{i}\;=\;S_{i}\,q$ for all $i\in\{1,\dots,n\}$, such that $E(\Gamma_{r})^{T}\begin{bmatrix}(S_{1}q)^{T}&(S_{2}q)^{T}&\cdots&(S_{n}q)^{T}\end{bmatrix}^{T}=0$. Note that $q\in\mathbb{R}^{2}$ has two degrees of freedom. Hence, any vector in $\mathrm{Null}(E(\Gamma_{r})^{T})=\mathrm{Null}(Q)$ lies in $\mathrm{IM}(V_{0})$, where
$V_{0}=\left[\begin{smallmatrix}S_{1}e_{1}&S_{1}e_{2}\\ \vdots&\vdots\\ S_{n}e_{1}&S_{n}e_{2}\end{smallmatrix}\right],$ (8)
with $e_{1}=\begin{bmatrix}1&0\end{bmatrix}^{T}$ and $e_{2}=\begin{bmatrix}0&1\end{bmatrix}^{T}$. Therefore, $\dim\mathrm{Null}(Q)=2$, and $\mathrm{rank}(Q)=2n-\dim\mathrm{Null}(Q)=2n-2$. ∎
We now examine the dynamics of the closed-loop system (7) to show that the proposed distributed control law drives the agents from any initial condition to the desired symmetric configuration. We derive the explicit solution of (7), showing that the limit configuration corresponds to the orthogonal projection of the initial state onto the subspace of $\mathcal{C}_{n}$-symmetric formations.
@@
Consider a MAS consisting of $n$ integrator agents (2), whose interaction topology is defined by $\mathcal{G}_{I}$, and let
@@
where $V_{0}$ is given in (8). Furthermore, the steady-state of each agent is given by
$\lim_{t\rightarrow\infty}p_{i}(t)=\frac{1}{n}S_{i}\sum_{k=1}^{n}S_{k}^{T}p_{k}(0).$ (10)